# tableReferences for Table S1

### v1

@@ A.3 Details about Wasserstein and Sliced-Wasserstein distances
A.5 Sliced-Wasserstein, MMD and C2ST as optimization target
@@
A.5.1 Fitting a Gaussian with gradient descent
@@
A.5.2 Fitting a mixture of Gaussians with Expectation-Maximisation
@@
E-step: For each of the $N$ datapoints $\hat{x}_{i}$ from $p_{\text{true}}$, we calculated the probability of it belonging mixture component 1 or 2:
@@
A.6 Additional scaling experiments with different sample size budgets and ranges
In Fig. 8, we evaluated the robustness of the measures against the number of samples and the dimensionality of the data. We observed notably poor performance of the measures in scenarios with limited data. Here, we further examine the performance of the distances across datasets of varying sample sizes, particularly for small sample set sizes, ranging from only 8 to 80 samples per set (Fig. S3). We examine three distinct data configurations where the distinction between the true and approximated distributions progressively increases from subpanels S3 a to c. Across all distances, it becomes evident that the larger the disparity between the two distributions, the fewer samples are needed for differentiation. In the experiment where all dimensions are mean-shifted by one, a sample size of 8 is sufficient to distinguish between the distributions. However, for less distinct distributions, such as the unimodal Gaussian or a mean-shift by one in only one dimension (Supp. Fig. S3 a, b), all distances exhibit poor performance in distinguishing between the distributions.
@@ A.6 Additional scaling experiments with different sample size budgets and ranges
A.7 Additional scaling experiments for different dimensionality of the data
@@
Figure S4: The impact of dimensionality varies across distances, with certain distances facing particular challenges in higher dimensional spaces. We compare sample sets of varying dimensionality (between 5 and 1000) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data. The bandwidth parameter in Gaussian Kernel MMD is set to 10 for all experiments. (a) Distances for a sample set from an n-dimensional standard normal distribution, for which the first dimension is shifted by one. (b) Distances for a sample set from an n-dimensional standard normal distribution, for which all dimensions are shifted by one. (c) Distances for a sample set from an n-dimensional standard normal distribution, for which variances are increased by one for all dimensions.
A.8 Sensitivity of the MMD bandwidth parameter
@@
Figure S5: The bandwidth parameter in Gaussian Kernel MMD is a sensitive parameter that requires careful selection for each dataset. The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data. (a) $MMD^{2}$ distance with varying bandwidth parameters between 0.1 and 5 for the 2d-MoG example compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) $MMD^{2}$ distance with varying bandwidth parameters between 0.1 and 20 for a 10-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) $MMD^{2}$ distance with varying bandwidth parameters between 1 and 40 for a 100-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
@@
Figure S6: Evaluating the effect of varying bandwidth parameters in Gaussian Kernel MMD for different sample sizes and dataset dimensionalities. (a,b) We compare sample sets with varying sample sizes (between 50 and 4k samples per set) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. (a) $MMD^{2}$ distance with varying bandwidth parameters (0.5, 1, 2) for the 2d-MoG example shown in Fig. 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) $MMD^{2}$ distance with varying bandwidth parameters (1, 3, 5) for a ten dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) $MMD^{2}$ distance with varying bandwidth parameters (5, 10, 40) based on 10k samples from a standard normal distributions with varying dimensions (between 5 and 1000). As in (b) the first dimension is shifted by one for the ’shifted’ dataset. Here we show one run due to computational costs.
A.9 Details about scientific application examples
@@
For the motion discrimination task, we used the decision times of a single animal during both correct and erroneous trials with dot motion coherence of 12.8%, leading to a one-dimensional dataset of 587 samples. DDMs were implemented using the pyDDM toolbox (Shinn et al., 2020). DDM1 used a linear drift and exponential decision boundaries. In contrast, DDM2 used a constant drift and a constant decision boundary. Both were sampled 1,000 times to create the two synthetic datasets. The real chest X-ray dataset consists of 70,153 samples, the generated datasets from PGGAN and SD consist of 10,000 and 2,352 samples respectively.

### v2

@@ A.4 Dependence of SW Distance on Number of Projections
A.6 Sample and computational complexity
Table S2: Summary of computational complexity and theoretical properties of metrics in terms of number of samples $N$ and data dimensionality $D$. Sample complexity here refers to the convergence rate of the sample-based estimate to the true value of the metric. *Bound based on Ghosal & Sen (2019), Nguyen & Ho (2024) for SWD and Gretton et al. (2012a) for MMD. Best case scenario. In practice, the computational cost of training a neural network scales superlinearly with both sample size and data dimensionality. General case (see Sec. 2.3, A.6 for details).Cost of calculating the square root of the covariance matrix in Eq. 4.
SW C2ST MMD FID
Sample Complexity ($N$) $\mathcal{O}(N^{-1/2})$* N/A $\mathcal{O}(N^{-1/2})$* N/A
Computational Complexity ($N$) $\mathcal{O}(N\log N)$ $\mathcal{O}(N)^{\dagger}$ $O(N^{2})^{\ddagger}$ $\mathcal{O}(N^{2})$
Computational Complexity ($D$) $\mathcal{O}(D)$ $\mathcal{O}(D)^{\dagger}$ $\mathcal{O}(D)$ $\mathcal{O}(D^{3})$
Estimator unbiased? no no yes no
We here highlight that the different distances also differ in their sample and computational complexities (Table S2). With respect to the number of samples $N$, MMD and FID have a complexity of $O(N^{2})$. Note that for MMD the computational cost can be reduced, e.g., by increasing variance or for a specific kernel choice (Gretton et al., 2012a, Zhao & Meng, 2015, Cheng & Xie, 2021, Bodenham & Kawahara, 2023, Bharti et al., 2023, Gretton et al., 2012b)). While SW distance scales with $O(N\log N)$ (Nadjahi, 2021), it is difficult to make principled assessments of the computational complexity for C2ST, as it is highly dependent on the chosen classifier. However, as more samples lead to larger training and test datasets, the sample size is likely to influence the compute time. Similarly, the computational complexity of computing these distances increases as the dimensionality of the data increases, with non-trivial scaling depending on the task and chosen hyperparameter. We also report the theoretical convergence of sample-based estimates for MMD, as well as SW, which is subject to active research. We report bounds from recent works (Ghosal & Sen, 2019, Gretton et al., 2012a). We do not report sample complexities for C2ST and FID, as these strongly depend on the choice of classifier and embedding network, respectively.
Note that despite their differences, all presented distances are reasonably tractable in the settings of our experiments, whereas the scaling experiments might be computationally unfeasible for other distances or datasets. Therefore, we strongly recommend carefully considering the complexity of the measure before conducting experiments on high-dimensional or very large datasets. We report the practical computation times for our experiments in Appendix A.11.
A.7 Mode coverage properties
Mode coverage is the ability of a model to capture and generate diverse data, i.e., from multiple modes of the underlying distribution (see Fig. S2;  S3). The community has focused on evaluation of mode coverage with different metrics driven by the development of GANs (Goodfellow et al., 2014, Gui et al., 2023, Saad et al., 2024). Empirically, in training generative models, mode coverage has been found to trade off with sample quality and speed, illustrating the generative learning trilemma (Xiao et al., 2022). All presented metrics can distinguish between a collapsed and a full distribution (Fig. 9), an empirical finding also reported in previous work (Che et al., 2017, Li et al., 2017, Deshpande et al., 2018, Borji, 2019). However, their sensitivity relies on different factors: SWD captures different modes when they are separated in one-dimensional projections, MMD depends on appropriate kernel choice, FID on expressive embeddings, and C2ST on well-trained classifiers. Other metrics used to quantify mode collapse include precision and recall (Kynkäänniemi et al., 2019).
A.8 Sliced-Wasserstein, MMD and C2ST as optimization target
@@
A.8.1 Fitting a Gaussian with gradient descent
@@
A.8.2 Fitting a mixture of Gaussians with Expectation-Maximisation
@@
E-step: For each of the $N$ datapoints $\hat{x}_{i}$ from $p_{\text{true}}$, we calculated the probability of it belonging to mixture component 1 or 2:
@@
A.9 Additional scaling experiments with different sample size budgets and ranges
In Fig. 8, we evaluated the robustness of the measures against the number of samples and the dimensionality of the data. We observed notably poor performance of the measures in scenarios with limited data. Here, we further examine the performance of the distances across datasets of varying sample sizes, particularly for small sample set sizes, ranging from only 8 to 80 samples per set (Fig. S4). We examine three distinct data configurations where the distinction between the true and approximated distributions progressively increases from subpanels S4 a to c. Across all distances, it becomes evident that the larger the disparity between the two distributions, the fewer samples are needed for differentiation. In the experiment where all dimensions are mean-shifted by one, a sample size of 8 is sufficient to distinguish between the distributions. However, for less distinct distributions, such as the unimodal Gaussian or a mean-shift by one in only one dimension (Supp. Fig. S4 a, b), all distances exhibit poor performance in distinguishing between the distributions.
@@ A.9 Additional scaling experiments with different sample size budgets and ranges
A.10 Additional scaling experiments for different dimensionality of the data
@@
Figure S5: The impact of dimensionality varies across distances, with certain distances facing particular challenges in higher dimensional spaces. We compare sample sets of varying dimensionality (between 5 and 1000) of a ‘true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data. The bandwidth parameter in Gaussian Kernel MMD is set to 10 for all experiments. (a) Distances for a sample set from an n-dimensional standard normal distribution, for which the first dimension is shifted by one. (b) Distances for a sample set from an n-dimensional standard normal distribution, for which all dimensions are shifted by one. (c) Distances for a sample set from an n-dimensional standard normal distribution, for which variances are increased by one for all dimensions.
A.11 Comparisons of practical compute times of different measures
Before computing such measures, in particular for scaling experiments such as the ones presented here, where measures are calculated across a large range of sample sizes $N$ and data dimensions $D$, the practical compute time of the chosen measures should be considered. Depending on the downstream application, it might be time-critical to quickly evaluate distances which might favor some measures over others. Aligned with theoretical considerations regarding sampling complexity etc. as presented in Table S2), empirical compute times vary between the different measures. Given that empirical computational times for a single measure itself vary depending on the exact implementation, compute infrastructure, and problem at hand, we list approximated compute times for running the scaling experiments in Table S3. The calculated runtime combines both the comparisons of the ‘true‘ and the ‘shifted‘ or ‘approximated‘ experiments. Each experiment contains five repeats across different sampled data subsets. The sample size experiment contains eight different sample size values $N$ (50, 100, 200, 500, 1000, 2000, 3000, 4000), and the dimensionality experiment scales tests six different $D$ (5, 10, 50, 100, 500, 1000). For more details, see Section 3. The version of C2ST we use here, which is based on NN-based classifiers, takes orders of magnitude longer to compute than SW and MMD. Note, however, that alternative implementations and classifier variants could speed this up.
Table S3: CPU wallclock run times for the comparison scaling experiments in Fig.8. The runtime combines both the comparisons of the true vs. the shifted or approximated distribution. Values are rounded estimates.
sample size experiment dimensionality experiment
2-dim MoGs 10-dim Gaussian n-dim Gaussian
SWD 0.3 s 0.3 s 1.5 s
C2ST 150 s 300 s 1500 s
MMD 3 s 3 s 80 s
A.12 Sensitivity of the MMD kernel choices and hyperparameters
The general formulation of MMD allows for a wide range of kernel choices, each potentially with their own hyperparameters. These choices can have significant effect on its behavior. We provide some experiments demonstrating of the importance of well-tuned bandwidth parameters for Gaussian Kernel MMD as well as the impact of different kernels.
@@ A.12 Sensitivity of the MMD kernel choices and hyperparameters
Figure S6: The bandwidth parameter in Gaussian Kernel MMD is a sensitive parameter that requires careful selection for each dataset. The sample size is fixed to $10$k for all experiments and we show the mean and standard deviation over five runs of randomly sampled data. (a) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters between 0.1 and 5 for the 2d-MoG example compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters between 0.1 and 20 for a 10-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters between 1 and 40 for a 100-dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example.
We first vary the bandwidth parameter with fixed sample sizes for the three example datasets used in Section 3 (Fig. S6). We show that the estimated $\mathsf{MMD}$ values vary significantly across bandwidths, and both setting the bandwidth too low or too high yield poor results. However, we note that the values yielded by the median heuristic (bandwidths of 1, 5, and 10 for the three datasets, respectively, as shown in the main text) are quite near the peaks of the curves at which MMD most effectively distinguishes the distributions.
@@
Figure S7: Evaluating the effect of varying bandwidth parameters in Gaussian Kernel MMD for different sample sizes and dataset dimensionalities. (a,b) We compare sample sets with varying sample sizes (between 50 and 4k samples per set) of a ’true’ distribution either with a second dataset of the same distribution or with a sample set from an approximated/shifted distribution. We show the mean and standard deviation over five runs of randomly sampled data. (a) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters (0.5, 1, 2) for the 2d-MoG example shown in Fig. 1 compared to samples from a unimodal Gaussian approximation with the same mean and covariance. (b) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters (1, 3, 5) for a ten dimensional standard normal distribution, for which the first dimension is shifted by one for the shifted example. (c) $\mathsf{MMD}^{2}$ distance with varying bandwidth parameters (5, 10, 40) based on 10k samples from a standard normal distributions with varying dimensions (between 5 and 1000). As in (b) the first dimension is shifted by one for the ’shifted’ dataset. Here we show one run due to computational costs.
We then choose a set of bandwidth parameters to compare across the scaling experiments of Section 3 (Fig. S7). Again, poor choices of bandwidth values give misleading results, but bandwidth choices guided by the median heuristic generally perform well.
Figure S8: Comparison of Gaussian Kernel MMD to different MMD kernels without tunable parameters. We compare the performance of $\mathsf{MMD}^{2}$ as presented in Fig.8 with a Gaussian kernel with bandwidth parameters adjusted for each dataset (1,5,10) (top row) to an MMD implementation with a linear kernel (middle row, $\mathsf{MMD}_{lin}$; $k(x,y)=\langle x,y\rangle$) and an energy-distance based kernel, i.e., the kernel induced by the Euclidean distance Sejdinovic et al. (2013) (bottom row, $\mathsf{MMD}_{en}$; $k(x,y)=\|x\rVert^{p}+\|x\rVert^{p}-\|x-y\rVert^{p}$, with $p=2$). The experiments, parameters for the Gaussian kernel bandwidth (indicated in the y-labels), and sample sizes etc. are identical to Fig. 8.
Finally, we vary the kernel choice across the scaling experiments of Section 3 (Fig. S7), using both a linear kernel ($\mathsf{MMD}_{lin}$) and the distance-induced kernel corresponding to the standard energy distance ($\mathsf{MMD}_{en}$). As expected, the linear kernel fails to distinguish distributions with matching means (2d-MoG) but performs reasonably well for distributions with mean-offsets even at high dimensions. The energy kernel performs similarly to the Gaussian kernel, without the added dependence on sensitive hyperparameters.
A.13 Additional results for ImageNet generative models
We generated 50,000 images using an unconditional diffusion model explicitly trained on ImageNet 64x64, as well as 100,000 class-conditional images from a conditional variant of the same model (Dockhorn et al., 2022) (we refer to it as GENIE). For additional comparison, we also generated 50,000 images using a consistency model explicitly trained on ImageNet 64x64 (Song et al., 2023) (we refer to it as CM). We compare these generated images to the ImageNet 64x64 test set.
For further comparison, we also evaluated image-generative models not specifically trained to reproduce ImageNet but designed for general-purpose image generation, such as Stable Diffusion and Midjourney (Rombach et al., 2022b, Midjourney, 2022). While these models might generate images that are more appealing to human observers, they do not necessarily produce images that align with the images contained in the ImageNet test set. We use the recently published million-scale dataset GenImage (Zhu et al., 2023). Especially, we include the models BigGAN (Brock et al., 2019), ablated diffusion model (ADM; Dhariwal & Nichol 2021), Glide (Nichol et al., 2022), Vector Quantized Diffusion Model (VQDM; Gu et al. 2022), Wukong (Wukong, 2022), Stable diffusion 1.5 ( SD1.5; Rombach et al. 2022b) and Midjourney  (Midjourney, 2022).
It’s important to note that not all models are specifically trained to capture ImageNet 64x64 images. For instance, models like Stable Diffusion and Midjourney are trained on much larger datasets (Schuhmann et al., 2022, Lin et al., 2014). Additionally, most of the models mentioned, except for ADM and BigGAN, are text-to-image generative models. To minimize significant distribution shifts, these models were prompted with the phrase "photo of [ImageNet class]" Zhu et al. (2023). Furthermore, all the other models generate images of larger resolution, which we resized to 64x64. Hence, we also only compare low-resolution features of natural images.
[…]
Overall, this analysis highlights that the choice of metric (i.e., what features it compares) and the specific implementation details (such as the classifier in C2ST estimates) matter and can lead to varying results.
Table S4: Evaluating discrepancy to the ImageNet test set.Each row presents various metrics computed on the Inception v3 embeddings of images. Columns correspond to different generative models. The initial three models are trained on ImageNet 64x64, serving as the reference point for comparison. Subsequent models are trained on alternative datasets or higher-resolution versions. In bold, we highlight the lowest value for each section of the table.
GENIE CM BigGAN ADM Glide VQDM WK SD1.5 Midjourney
FID $\mathbf{5.1\cdot 10^{0}}$ $5.3\cdot 10^{0}$ $1.2\cdot 10^{1}$ $1.1\cdot 10^{1}$ $1.1\cdot 10^{1}$ $\mathbf{9.8\cdot 10^{0}}$ $1.1\cdot 10^{1}$ $1.2\cdot 10^{1}$ $1.0\cdot 10^{1}$
SWD $22.3\cdot 10^{-}$ 2 $\mathbf{2.2\cdot 10^{-}2}$ $25.3\cdot 10^{-}$ 2 $25.1\cdot 10^{-}$ 2 $24.9\cdot 10^{-}$ 2 $\mathbf{4.0\cdot 10^{-}2}$ $24.8\cdot 10^{-}$ 2 $25.0\cdot 10^{-}$ 2 $24.5\cdot 10^{-}$ 2
[…]
A.14 Details about scientific application examples
@@
For the motion discrimination task, we used the decision times of a single animal during both correct and erroneous trials with dot motion coherence of 12.8%, leading to a one-dimensional dataset of 1023 samples. From these 80% were used as a train set and 20% as a test set DDMs were implemented using the pyDDM toolbox (Shinn et al., 2020). DDM1 used a linear drift and exponential decision boundaries. In contrast, DDM2 used a constant drift and a constant decision boundary. Both were sampled 1,000 times to create the two synthetic datasets. The real chest X-ray dataset consists of 70,153 train samples and 25,596 test samples, the generated datasets from PGGAN and SD consist of 10,000 and 2,352 samples respectively.